Burt Rosenberg

Problem Set 4 _{Due: 18 October, 1994}^{Out: 11 October, 1994}

Goals

Understanding recursion.

Reading Assignment

Read chapter 16 Condensed Pascal. For the programming assignment, you will also need pages 118–119 in chapter 5.

Programming Assignment

Exercises 16–1, 16–2, 16–3, page 367 of Condensed Pascal. Extra Credit: Exercise 16–8.

Midterm Announcement

In class midterm, Thursday, 20 October. The midterm will be a 75 minute closed-book written exam on all material up to recursion.